OpenStudy (anonymous):

how to solve this one? solve for x: (x^2)(2^x)-2(x)=0

OpenStudy (aum):

\[ (x^2)(2^x)-2(x)=0 \\ x(x2^x-2) = 0 \\ x = 0 ~~\text{ or }~~ x2^x = 2 \]Just by inspection, x = 1 works for the second equation. x = 0 or x = 1.